# Relative monodromy of ramified sections on abelian schemes

### v3

@@ 1 Introduction
2 Monodromy on abelian schemes
2.1 Setting
In general if $X$ is a quasi-projective complex algebraic variety, with $X(\mathbb{C})$ we denote the set of its closed points which has a structure of complex manifold.
Let $S$ be a regular, irreducible, quasi-projective variety defined over $\mathbb{C}$; we also tacitly assume that $S$ is maximal. Let $\phi:\mathcal{A}\to S$ be an abelian scheme over $S$ of relative dimension $g\geq 1$, namely a proper smooth group $S$-scheme admitting a “zero section” $\sigma_{0}:S\to\mathcal{A}$ and whose fibers are $g$ dimensional abelian varieties. For any $s\in S(\mathbb{C})$ the fiber over $s$ is denoted by $\mathcal{A}_{s}$. We denote with $\Sigma^{\operatorname{an}}$ the sheaf over $S(\mathbb{C})$ of holomorphic sections of $\phi$ whereas $\Sigma(S)\subset\Sigma^{\operatorname{an}}(S(\mathbb{C}))$ is the abelian group of (global) algebraic $\mathbb{C}$-sections.
In Section 3.1 we will show that our main results are insensitive to finite base changes of $S$ and to isogenies. Hence, by using for instance [13, Proof of Theorem B.1 Devissages $(iv)-(vi)$] we can assume that the abelian scheme carries principal polarization and it has $-\ell$-structure.
In this paper, sometimes we also need to assume an analytic viewpoint on families of abelian varieties: we look at $S(\mathbb{C})$ and $\mathcal{A}(\mathbb{C})$ as normal quasi-projective varieties with a surjective holomorphic map $\phi:\mathcal{A}(\mathbb{C})\to S(\mathbb{C})$ whose fibers are abelian varieties, thus the zero section $\sigma_{0}:S(\mathbb{C})\to\mathcal{A}(\mathbb{C})$ and the fiberwise group operations are complex analytic. The previous family can be regarded as a geometric model of a polarized abelian variety $A$ defined over the function field $\mathbb{C}(S)$ of meromorphic functions on $S(\mathbb{C})$: from the latter perspective a rational point $\sigma$ of $A$ over $\mathbb{C}(S)$ is equivalently an algebraic section of $\phi:\mathcal{A}(\mathbb{C})\to S(\mathbb{C})$.
Fix an integer $\ell\geq 3$. Now we use two points of view to describe the moduli space of $g$-dimensional principally polarized abelian varieties, with $-\ell$-structure. First we use scheme theory. There exist an irreducible smooth quasi-projective variety $\mathbb{A}_{g}$ over $\overline{\mathbb{Q}}$ and a principally polarized abelian scheme $\mathfrak{A}_{g}\to\mathbb{A}_{g}$ of relative dimension $g$ with symplectic $-\ell$-structure with the following property: if $S$ is any scheme over $\mathbb{C}$ and $\mathcal{A}\to S$ is a principally polarized abelian scheme of relative dimension $g$ with symplectic $-\ell$-structure, then there exists a unique $\mathbb{C}$-morphism $p:S\to\mathbb{A}_{g}$ such that $\mathcal{A}$ is isomorphic to the pull-back $\mathfrak{A}_{g}\times_{\mathbb{A}_{g}}S$ (i.e. $\mathbb{A}_{g}$ is a fine moduli space). Let $\pi:\mathfrak{A}_{g}\to\mathbb{A}_{g}$ be the universal abelian variety and write $p_{\mathcal{A}}$ for the induced $\mathbb{C}$-morphism $\mathcal{A}\to\mathfrak{A}_{g}$. Also $\mathfrak{A}_{g}$ is an irreducible, smooth, quasi-projective variety definable over a number field. In other words, we are supposed to have the following commutative diagram:
@@ 2.1 Setting
Now, let’s assume a more analytic viewpoint on moduli spaces introducing modular families of abelian varieties. We denote by $\mathbb{H}_{g}$ the Siegel upper-half plane:
$\mathbb{H}_{g}:=\{\tau\in\textnormal{Sym}(g,\mathbb{C})\mid\textnormal{Im}\,\tau>0\}.$
The isomorphism classes of principally polarized abelian varieties are in one-to-one correspondence with $\mathbb{H}_{g}/\textnormal{Sp}_{2g}(\mathbb{Z})$ where $\textnormal{Sp}_{2g}(\mathbb{Z})$ is a discrete subgroup of $\textnormal{Sp}_{2g}(\mathbb{R})$. To any torsion-free subgroup $\Gamma\subset\textnormal{Sp}_{2g}(\mathbb{Z})$ of finite index one can associate a universal family $\pi_{\Gamma}:\mathcal{A}_{\Gamma}(\mathbb{C})\to S_{\Gamma}(\mathbb{C})$ of principally polarized abelian varieties: following Satake we call such a family $\pi_{\Gamma}$ a Kuga family of principally polarized abelian varieties. Without loss of generality, we always assume that $\Gamma$ is torsion-free. Given a principally polarized abelian scheme $\phi:\mathcal{A}\to S$ of relative dimension $g$ with level $\ell$-structure, there is a congruence subgroup
$\Gamma=\Gamma(\ell)\subset\textnormal{Sp}_{2g}(\mathbb{R})$
and a unique modular map $p:S\to\mathbb{H}_{g}/\Gamma=S_{\Gamma}$ such that $\phi:\mathcal{A}(\mathbb{C})\to S(\mathbb{C})$ is isomorphic to the pull-back of the Kuga family $\pi:\mathcal{A}_{\Gamma}(\mathbb{C})\to S_{\Gamma}(\mathbb{C})$ by $p$. For our purposes we are interested in some vanishing results of the Mordell-Weil rank. Now, we fix once and for all a suitable level $\ell\geq 3$ such that the Kuga family $\overline{\pi}_{\Gamma}:\overline{\mathcal{A}}_{\Gamma}(\mathbb{C})\to\overline{S}_{\Gamma}(\mathbb{C})$ has no locally constant parts, i.e. there is no constant part and the same is true even when $\Gamma$ is replaced by any subgroup of finite index.
@@ 2.2 Periods and abelian logarithms
Let us consider an abelian scheme $\phi:\mathcal{A}\to S$ with a zero-section as above. Each fiber $\mathcal{A}_{s}(\mathbb{C})$ is analytically isomorphic to a complex torus $\mathbb{C}^{g}/\Lambda_{s}$ and for any subset $W\subseteq S(\mathbb{C})$ we denote $\Lambda_{W}:=\bigsqcup_{s\in W}\Lambda_{s}$. The space $\operatorname{Lie}(\mathcal{A}):=\bigsqcup_{s\in S(\mathbb{C})}\operatorname{Lie}(\mathcal{A}_{s})$ has a natural structure of $g$-dimensional holomorphic vector bundle $f\colon\operatorname{Lie}(\mathcal{A})\to S(\mathbb{C})$ (it is actually a complex Lie algebra bundle). By using the fiberwise exponential maps one can define a global map $\exp\colon\operatorname{Lie}(\mathcal{A})\to\mathcal{A}$. Let $\Theta_{0}\subset\mathcal{A}$ be the image of the zero section of the abelian scheme, then obviously
@@
Clearly $S(\mathbb{C})$ can be covered by finitely many open simply connected subsets where the holomorphic vector bundle $f\colon\operatorname{Lie}(\mathcal{A})\to S(\mathbb{C})$ trivializes. Let $U\subseteq S(\mathbb{C})$ be any of such subsets and consider the induced holomorphic map $f\colon\Lambda_{U}\to U$; it is actually a fiber bundle with structure group $\textnormal{GL}(n,\mathbb{Z})$. Since $U$ is simply connected, by [12, Lemma 4.7] we conclude that $f\colon\Lambda_{U}\to U$ is a topologically trivial fiber bundle. Thus we can find $2g$ continuous sections of $f$:
@@
2.3 Monodromy of periods
@@
Let $X$ be a locally contractible topological space and denote by $\mathbb{Z}_{X}$ the constant sheaf of abelian groups over $X$ with constant stalk $\mathbb{Z}$. Then we have a canonical isomorphism
$H^{q}_{\textrm{sing}}(X,\mathbb{Z})\cong H^{q}(X,\mathbb{Z}_{X}),$
where we are considering the cohomology of $X$ with coefficients in the constant sheaf of stalk $\mathbb{Z}_{X}$ on the right, and the singular cohomology with coefficients in $\mathbb{Z}$ on the left. The same result holds with $\mathbb{Z}$ replaced by any commutative ring $G$.
Now, let $\phi:\mathcal{A}\to S$ be an abelian scheme as defined above and consider the constant sheaf $\mathbb{Z}_{\mathcal{A}}$. We want to define a sheaf on the base $S$ containing information on periods of the abelian schemes, to this end let us consider the direct image functor $\phi_{*}$. Since $\phi_{*}$ is left exact and the category of sheaves of abelian groups has enough injectivities, the right derived functors of $\phi_{*}$ are well defined from the category of sheaves on $\mathcal{A}$ to the category of sheaves on $S$: they are called higher direct image functors and will be denoted by $R^{k}\phi_{*}$. For any $k\geq 0$, define the sheaf
$\mathcal{P}_{S}^{k}:=R^{k}\phi_{*}\mathbb{Z}_{\mathcal{A}}.$
By [21, Proposition 8.1], this is the sheaf on $S$ associated to the presheaf
$U\mapsto H^{k}(\phi^{-1}(U),\mathbb{Z}_{\mathcal{A}}|_{\phi^{-1}(U)}).$
Since $\phi:\mathcal{A}\to S$ is a smooth surjective morphism of algebraic varieties over $\mathbb{C}$, by [21, Theorem 10.4] then $\phi$ is a submersion. We can apply Ehresmann’s Lemma to conclude that the proper submersion $\phi$ is a $C^{\infty}$-fiber bundle.
Since $S(\mathbb{C})$ is locally contractible, then for sufficiently small $U\subseteq S(\mathbb{C})$, the open sets $\mathcal{A}_{U}:=\phi^{-1}(U)$ have the same homotopy type as the fibre $\mathcal{A}_{s}(\mathbb{C})$ with $s\in U$. Using the invariance under homotopy, i.e. the fact that if $U$ is a contractible space then $H^{k}(\mathcal{A}_{s}(\mathbb{C})\times U,\mathbb{Z})=H^{k}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})$ for all $k\geq 0$, we deduce that the sheaves $\mathcal{P}_{S}^{k}=R^{k}p_{*}\mathbb{Z}_{\mathcal{A}}$ are locally constant sheaves (or equivalenty local systems) on $S$ with stalk
$(\mathcal{P}_{S}^{k})_{s}=(R^{k}\phi_{*}\mathbb{Z}_{\mathcal{A}})_{s}=H^{k}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z}).$
@@
The homeomorphism $\psi$ induces a group automorphism $\psi^{k}$ of $H^{k}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})$, for any $k\geq 0$. Therefore, we obtain the monodromy representations
$\rho^{k}:\pi_{1}(S(\mathbb{C}),s)\to\textrm{Aut}\,H^{k}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})$ (7)
@@
$\rho^{k}(\gamma)(\eta):=\psi^{k}\eta,\qquad\eta\in H^{k}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z}).$
Thus, thanks to the fact the local systems we are considering is obtained using a fibration, the monodromy representation is in fact induced by homeomorphisms of the fibre $\mathcal{A}_{s}(\mathbb{C})$ onto itself. In particular, it follows that the monodromy representation on the cohomology of the fibre $H^{\bullet}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})$ is compatible with the cup-product on $H^{\bullet}(\mathbb{A}_{s}(\mathbb{C}),\mathbb{Z})$, in the sense that each $\rho(\gamma)\in\textrm{Aut}\,H^{\bullet}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})$ is a ring automorphism for the ring structure given by the cup-product.
Recall that by Poincaré duality we have isomorphisms $H^{k}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})\cong H_{2g-k}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})$. Moreover, by the Künneth formula the groups $H_{k}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})$ and $H^{k}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})$ are free abelian groups of finite rank $\binom{2g}{k}$ for all $k=1,\ldots,2g$. Therefore, we can look at $\mathcal{P}_{S}^{2g-1}$ as a local system with constant stalk $\mathbb{Z}^{2g}$ to be identified with singular homology; we will simply write $\mathcal{P}$ instead of $\mathcal{P}_{S}^{2g-1}$. In other words, from now on we fix $k=2g-1$ and we only deal with the monodromy action induced on the first homology group identifying $\textrm{Aut}\,H_{1}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})\cong\textrm{GL}_{2g}(\mathbb{Z})$:
@@
Thus, monodromy representation of periods introduced in Equation 7 can be actually though as a map:
@@
Since $S(\mathbb{C})$ is path connected, all the groups $\textrm{Mon}(\mathcal{A},s)$ are conjugate when we vary the base point $s\in S(\mathbb{C})$: in other words, the monodromy group is defined up to an inner automorphism of the group $\textrm{Sp}_{2g}(\mathbb{Z})$. Thus, fix once and for all a base point $s\in S(\mathbb{C})$ and denote the group $\textrm{Mon}(\mathcal{A},s)$ simply by $\textrm{Mon}(\mathcal{A})$, without writing any dependencies on the base point.
@@
Periods from universal family.
In the case of a universal family, a more explicit description of the monodromy action on periods is described in [17, Section 3]. Keeping the same notations as above, let us denote by $\pi:\mathfrak{A}_{g}\to\mathbb{A}_{g}$ the universal family of principally polarized $g$-dimensional abelian varieties with $-\ell$-structure for some $g\geq 1$ and $\ell\geq 3$. Let $\mathbb{H}_{g}$ denote Siegel’s upper half space, i.e., the symmetric matrices in $\textrm{Mat}_{g\times g}(\mathbb{C})$ with positive definite imaginary part. We have holomorphic uniformizing maps

### v4

@@ 1 Introduction
2 Generalities on abelian schemes
2.1 Ramification and torsion
We discuss here the notion of unramified section formally given in LABEL:{def:ram_sect}, providing examples and properties. We prove a characterization useful for constructing unramified sections, and we show that torsion sections are always unramified. We refer to the notation introduced at the beginning and to the diagram
@@ 2.1 Ramification and torsion
where $q^{\prime}$ is unramified, $q$ is dominant, $S^{\operatorname{ur}}$ is normal and $\nu$ is the normalization map.
In general, the normalization map $\nu\colon T^{\nu}\to T$ is not unramified, and understanding its ramification is an interesting problem. For instance, the normalization of a nodal curve is unramified but not étale, while the normalization of a cuspidal curve is ramified. On the other hand, if we start with a normal irreducible variety, the relative normalizations within finite field extensions of the function field yield natural examples of unramified morphisms. In fact, there is a strong classification result: étale covers of a normal integral variety, such as $T^{\nu}$, correspond exactly to the relative normalizations of $T^{\nu}$. A more general version of this statement, which includes non-integral schemes, can be found in [26, Exposé I, Corollaire 10.2 and Corollaire 10.3].
We aim to establish criteria for identifying unramified sections $\sigma\colon S\to\mathcal{A}$. At first glance, one might expect that unramified sections should be defined as those induced by unramified covers of $T$, rather than $T^{\nu}$. However, since $T$ may fail to be normal, this definition proves to be ill-behaved, as previously discussed. Nevertheless, we observe that due to the “cancellation property” of unramified morphisms (see [39, TAG 02GG]), sections arising from unramified covers of $T$ remain unramified under our current definition. The converse, however, does not hold in general, as the normalization map $\nu\colon T^{\nu}\to T$ may itself be ramified.
We remark that the normality condition is essential in order to have the correspondence between the notion of topological covers and finite étale covers. In fact while finite étale morphisms induce topological covers of the corresponding complex spaces, we need the normality condition to conclude that an unramified (finite) topological cover induces a finite étale map. This implies the first simple criterion to construct unramified sections:
Proposition 2.1.
Assume the modular map $p:S\to T$ to be finite. Let $\sigma:S\to\mathcal{A}$ be a section. Let $\mathcal{A}^{\prime\prime}/S^{\prime\prime}$ an abelian scheme with a cartesian diagram
[…]
where $f$ is an unramified morphism. If $\sigma$ is pullback of a section of the abelian scheme $\mathcal{A}^{\prime\prime}/S^{\prime\prime}$ by $g$, then $\sigma$ is an unramified section.
Proof.
By [26, Exposé I, Corollaire 9.11], the morphism $f$ is étale, and $S^{\prime\prime}$ is normal and irreducible. By [26, Exposé I, Corollaire 10.2 and Corollaire 10.3], it follows that $\sigma$ is the pullback of a section of the abelian scheme $\mathcal{A}^{\prime}/S^{\operatorname{ur}}$, thus it is unramified.
Let $A$ and $A^{\prime}$ be the generic fibres of $\mathcal{A}$ and $\mathcal{A}^{\prime}$ respectively, since the induced map $A\to A^{\prime}$ is dominant, any section $\sigma:S\to\mathcal{A}$ induces a $\mathbb{C}(S)$ point of $A^{\prime}$ that we denote by $\sigma^{\prime}$. The closure of $\sigma^{\prime}$ in $\mathcal{A}^{\prime}$ is denoted by $\Sigma^{\prime}$ and clearly $\mathbb{C}(\Sigma^{\prime})$ embeds into $\mathbb{C}(S)$. Moreover recall that $\phi^{\prime}\colon\Sigma^{\prime}\to S^{\operatorname{ur}}$ is surjective.
Proposition 2.2.
Keep the notations fixed above and in Equation 3. A section $\sigma\colon S\to\mathcal{A}$ is unramified if and only if there exists a locally closed subscheme $\mathcal{B}/S^{\operatorname{ur}}$ of $\mathcal{A}^{\prime}/S^{\operatorname{ur}}$ with the following property: the restriction $\phi^{\prime}:\mathcal{B}\to S^{\operatorname{ur}}$ is finite unramified. If $\sigma:S\to\mathcal{A}$ is a section such that $\Sigma^{\prime}\subseteq\mathcal{B}$, then $\sigma$ is unramified.
Proof.
Assume that $\sigma$ is unramified, so $\sigma=q^{\ast}\sigma^{\prime}$ where $\sigma^{\prime}\colon S^{\operatorname{ur}}\to\mathcal{A}^{\prime}$ is a section of $\phi^{\prime}$. Then $\Sigma^{\prime}=\sigma^{\prime}(S^{\operatorname{ur}})$ and it is enough to take $\mathcal{B}:=\Sigma^{\prime}$. Vice versa assume the existence of $\mathcal{B}$; by [39, Tag 02IK] $\mathcal{B}$ is a Noetherian scheme (so a Noetherian topological space) then it has a finite number of irreducible components. This implies that all the connected components are closed and open. Therefore, the restriction of the morphism $\phi^{\prime}\colon\mathcal{B}\to S^{\operatorname{ur}}$ to its connected component $\mathcal{B}^{\prime}$ that contains $\Sigma^{\prime}$ is finite and unramified. By [39, Tag 0GS9] (here we use the hypothesis that $S^{\operatorname{ur}}$ is normal) we conclude that $\phi^{\prime}:\mathcal{B}^{\prime}\to S^{\operatorname{ur}}$ is finite étale and $\mathcal{B}^{\prime}$ is irreducible. So $\mathcal{B}^{\prime}=\Sigma^{\prime}$ because they have the same dimension. We have shown that $\Sigma^{\prime}$ is a connected component of $\mathcal{B}$. Now since $\mathbb{C}(\Sigma^{\prime})\subseteq\mathbb{C}(S)$, and $\phi^{\prime}:\Sigma^{\prime}\to S^{\operatorname{ur}}$ is in particular unramified it follows that $\mathbb{C}(S^{\operatorname{ur}})=\mathbb{C}(\Sigma^{\prime})$, this because $\mathbb{C}(S^{\operatorname{ur}})$ is the maximal unramified extension of $\mathbb{C}(T^{\nu})$ in $\mathbb{C}(S)$. It follows that the morphism $\phi^{\prime}\colon\Sigma^{\prime}\to S^{\operatorname{ur}}$ is finite étale of degree $1$, hence an isomorphism. By [39, Tag 024U] there exists a section $\widetilde{\sigma}\colon S^{\operatorname{ur}}\to\mathcal{B}$ such that $\widetilde{\sigma}(S^{\operatorname{ur}})=\Sigma^{\prime}$. In particular $\widetilde{\sigma}$ is a section of $\phi^{\prime}:\mathcal{A}^{\prime}\to S^{\operatorname{ur}}$ and $q^{\ast}\widetilde{\sigma}=\sigma$. ∎
Corollary 2.3.
Let $m\in\mathbb{Z}_{>0}$ be a positive integer. A section $\sigma:S\to\mathcal{A}$ is unramified if and only if $m\sigma$ is unramified. In particular, any torsion section is unramified
Proof.
If $\sigma$ is unramified then $\sigma=q^{\ast}\sigma^{\prime}$ for $\sigma^{\prime}\colon S^{\operatorname{ur}}\to\mathcal{A}^{\prime}$, so $m\sigma=q^{\ast}(m\sigma^{\prime})$. Vice versa, denote by $f:\mathcal{A}\to\mathcal{A}^{\prime}$ the upper left horizontal map in Equation 3. Since $m\sigma$ is unramified, there exists a section $\psi:S^{\operatorname{ur}}\to\mathcal{A}^{\prime}$ such that $m\sigma=q^{\ast}\psi$. This condition implies that $[m]\circ f\circ\sigma=\psi\circ q$. Define $\Sigma^{\prime}:=f\circ\sigma(S)\subseteq\mathcal{A}^{\prime}$ and observe this corresponds to take the closure of the image of $\sigma$ in the generic fiber as above. Thus, $\Sigma^{\prime}$ is irreducible and clearly $\mathbb{C}(\Sigma^{\prime})$ embeds into $\mathbb{C}(S)$. The multiplication map $[m]:\mathcal{A}^{\prime}\to\mathcal{A}^{\prime}$ is finite étale by [10, V, Proposition 20.7]. Let us consider the locally closed subscheme $\mathcal{B}/S^{\operatorname{ur}}$ satisfying the following cartesian diagram:
[…]
Observe that the map $\mathcal{B}\to S^{\operatorname{ur}}$ is étale, because it is base change of an étale morphism. Since $\Sigma^{\prime}\subseteq\mathcal{B}$, the conclusion follows by 2.2. ∎
We now provide an example of unramified non-torsion section.
Example 2.4.
Let us denote by $\mathcal{L}\to S_{0}$ the Legendre elliptic scheme, where $S_{0}:=\mathbb{P}_{1}\setminus\{0,1,\infty\}$. Consider the product
$(\phi_{1},\phi_{2}):\mathcal{L}\times\mathcal{L}\to S_{0}\times S_{0},$
where we use the coordinates $(x,y)$ in $S_{0}\times S_{0}$. We let $X$ be the line in $S_{0}^{2}$ defined by $x+y=2$, so this line is isomorphic under the first projection to $B:=\mathbb{P}_{1}\setminus\{0,1,2,\infty\}$. By projection from $X$ to $B$, we may define a scheme $\phi:\mathcal{A}\to B$ over $B$ whose fibers are of the type $\phi^{-1}(\lambda)=\mathcal{L}_{\lambda}\times\mathcal{L}_{2-\lambda}$.
The abelian scheme $\phi:\mathcal{A}\to B$ may be interpreted as the fiber product of two elliptic schemes $\mathcal{E}_{1}\times_{B}\mathcal{E}_{2}\to B$, where $\mathcal{E}_{1}\to B$ is the Legendre scheme associating $\mathcal{L}_{\lambda}\to\lambda$ and $\mathcal{E}_{2}\to B$ is the Legendre scheme associating $\mathcal{L}_{\lambda}\to 2-\lambda$. Define the multisection:
$\tau(\lambda):=\left(\left(2,\sqrt{2(2-\lambda)}\right),\left(2,\sqrt{2\lambda}\right)\right).$
It becomes a non-torsion section on the base change induced by the degree $4$ field extension $\mathbb{C}(B)\subseteq\mathbb{C}(B)(\psi,\mu)$ where $\psi^{2}+\mu^{2}=2$ and $\mu^{2}=\lambda$. Such a field extension induces an unramified morphism onto $B$ and therefore this gives rise to an unramified (finite flat) modular map. This implies that $\tau$ is an unramified section.
@@ 2.2 Periods and abelian logarithms
Let us consider an abelian scheme $\phi:\mathcal{A}\to S$ with a zero-section as above. Each fiber $\mathcal{A}_{s}(\mathbb{C})$ is analytically isomorphic to a complex torus $\mathbb{C}^{g}/\Lambda_{s}$ and for any subset $W\subseteq S(\mathbb{C})$ we denote $\Lambda_{W}:=\bigsqcup_{s\in W}\Lambda_{s}$. The space $\operatorname{Lie}(\mathcal{A}):=\bigsqcup_{s\in S(\mathbb{C})}\operatorname{Lie}(\mathcal{A}_{s}(\mathbb{C}))$ has a natural structure of $g$-dimensional holomorphic vector bundle $f\colon\operatorname{Lie}(\mathcal{A})\to S(\mathbb{C})$ (it is actually a complex Lie algebra bundle). By using the fiberwise exponential maps one can define a global map $\exp\colon\operatorname{Lie}(\mathcal{A})\to\mathcal{A}(\mathbb{C})$. Let $\Theta_{0}\subset\mathcal{A}(\mathbb{C})$ be the image of the zero section of the abelian scheme, then obviously
@@
Clearly $S(\mathbb{C})$ can be covered by open simply connected subsets where the holomorphic vector bundle $f\colon\operatorname{Lie}(\mathcal{A})\to S(\mathbb{C})$ trivializes. Let $U\subseteq S(\mathbb{C})$ be any of such subsets and consider the induced holomorphic map $f\colon\Lambda_{U}\to U$; it is actually a fiber bundle with structure group $\textnormal{GL}(n,\mathbb{Z})$. Since $U$ is simply connected, by [15, Lemma 4.7] we conclude that $f\colon\Lambda_{U}\to U$ is a topologically trivial fiber bundle. Thus we can find $2g$ continuous sections of $f$:
@@
3 Monodromy groups
3.1 Monodromy of abelian schemes
@@
Let $X$ be a locally contractible topological space and denote by $\underline{\mathbb{Z}}$ the constant sheaf of abelian groups over $X$ with constant stalk $\mathbb{Z}$. Then we have a canonical isomorphism
$H^{q}_{\textrm{sing}}(X,\mathbb{Z})\cong H^{q}(X,\underline{\mathbb{Z}}),$
where we are considering the cohomology of $X$ with coefficients in the constant sheaf of stalk $\underline{\mathbb{Z}}$ on the right, and the singular cohomology with coefficients in $\mathbb{Z}$ on the left. The same result holds with $\mathbb{Z}$ replaced by any commutative ring $G$.
Now, let $\phi:\mathcal{A}\to S$ be an abelian scheme as defined above. We want to define a sheaf on the base $S$ containing information on periods of the abelian schemes, to this end let us consider the direct image functor $\phi_{*}$. Since $\phi_{*}$ is left exact and the category of sheaves of abelian groups has enough injectivities, the right derived functors of $\phi_{*}$ are well defined from the category of sheaves on $\mathcal{A}$ to the category of sheaves on $S$: they are called higher direct image functors and will be denoted by $R^{k}\phi_{*}$. For any $k\geq 0$, define the sheaf
$\mathcal{P}_{S}^{k}:=R^{k}\phi_{*}\underline{\mathbb{Z}}.$
By [28, Proposition 8.1], this is the sheaf on $S$ associated to the presheaf
$U\mapsto H^{k}(\phi^{-1}(U),\underline{\mathbb{Z}}_{|\phi^{-1}(U)}).$
Since $\phi:\mathcal{A}\to S$ is a smooth surjective morphism of algebraic varieties over $\mathbb{C}$, by [28, Theorem 10.4] then $\phi$ is a submersion. We can apply Ehresmann’s Lemma to conclude that the proper submersion $\phi$ is a $C^{\infty}$-fiber bundle.
Since $S(\mathbb{C})$ is locally contractible, then for sufficiently small $U\subseteq S(\mathbb{C})$, the open sets $\mathcal{A}_{U}:=\phi^{-1}(U)$ have the same homotopy type as the fibre $\mathcal{A}_{s}(\mathbb{C})$ with $s\in U$. Using the invariance under homotopy, i.e. the fact that if $U$ is a contractible space then $H^{k}(\mathcal{A}_{s}(\mathbb{C})\times U,\underline{\mathbb{Z}})=H^{k}(\mathcal{A}_{s}(\mathbb{C}),\underline{\mathbb{Z}})$ for all $k\geq 0$, we deduce that the sheaves $\mathcal{P}_{S}^{k}=R^{k}\phi_{*}\underline{\mathbb{Z}}$ are locally constant sheaves (or equivalenty local systems) on $S$ with stalk
$(\mathcal{P}_{S}^{k})_{s}=(R^{k}\phi_{*}\underline{\mathbb{Z}})_{s}=H^{k}(\mathcal{A}_{s}(\mathbb{C}),\underline{\mathbb{Z}}).$
@@
The homeomorphism $\psi$ induces a group automorphism $\psi^{k}$ of $H^{k}(\mathcal{A}_{s}(\mathbb{C}),\underline{\mathbb{Z}})$, for any $k\geq 0$. Therefore, we obtain the monodromy representations
$\rho^{k}:\pi_{1}(S(\mathbb{C}),s)\to\textrm{Aut}\,H^{k}(\mathcal{A}_{s}(\mathbb{C}),\underline{\mathbb{Z}})$ (9)
@@
$\rho^{k}(\gamma)(\eta):=\psi^{k}\eta,\qquad\eta\in H^{k}(\mathcal{A}_{s}(\mathbb{C}),\underline{\mathbb{Z}}).$
Thus, thanks to the fact the local systems we are considering is obtained using a fibration, the monodromy representation is in fact induced by homeomorphisms of the fibre $\mathcal{A}_{s}(\mathbb{C})$ onto itself. In particular, it follows that the monodromy representation on the cohomology of the fibre $H^{\bullet}(\mathcal{A}_{s}(\mathbb{C}),\underline{\mathbb{Z}})$ is compatible with the cup-product on $H^{\bullet}(\mathcal{A}_{s}(\mathbb{C}),\underline{\mathbb{Z}})$, in the sense that each $\rho(\gamma)\in\textrm{Aut}\,H^{\bullet}(\mathcal{A}_{s}(\mathbb{C}),\underline{\mathbb{Z}})$ is a ring automorphism for the ring structure given by the cup-product.
Recall that by Poincaré duality we have isomorphisms $H^{k}(\mathcal{A}_{s}(\mathbb{C}),\underline{\mathbb{Z}})\cong H_{2g-k}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})$. Moreover, by the Künneth formula the groups $H_{k}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})$ and $H^{k}(\mathcal{A}_{s}(\mathbb{C}),\underline{\mathbb{Z}})$ are free abelian groups of finite rank $\binom{2g}{k}$ for all $k=1,\ldots,2g$. Therefore, we can look at $\mathcal{P}_{S}^{2g-1}$ as a local system with constant stalk $\mathbb{Z}^{2g}$ to be identified with singular homology; we will simply write $\mathcal{P}$ instead of $\mathcal{P}_{S}^{2g-1}$. In other words, from now on we fix $k=2g-1$ and we only deal with the monodromy action induced on the first homology group identifying $\textrm{Aut}\,H_{1}(\mathcal{A}_{s}(\mathbb{C}),\mathbb{Z})\cong\textrm{GL}_{2g}(\mathbb{Z})$:
@@
Thus, monodromy representation of periods introduced in Equation 9 can be actually thought as a map:
@@
Since $S(\mathbb{C})$ is path connected, all the groups $\textrm{Mon}(\mathcal{A},s)$ are conjugate when we vary the base point $s\in S(\mathbb{C})$: in other words, the monodromy group is defined up to an inner automorphism of the group $\textrm{GL}_{2g}(\mathbb{Z})$. Thus, fix once and for all a base point $s\in S(\mathbb{C})$ and denote the group $\textrm{Mon}(\mathcal{A},s)$ simply by $\textrm{Mon}(\mathcal{A})$, without writing any dependencies on the base point.
@@
3.2 The case of universal families
For universal families, a more explicit description of the monodromy action on periods is described in [22, Section 3]. Let $\mathbb{H}_{g}$ denote Siegel’s upper half space, i.e., the symmetric matrices in $\textrm{Mat}_{g\times g}(\mathbb{C})$ with positive definite imaginary part. We have holomorphic uniformizing maps